# Multimodal Fusion Method with Spatiotemporal Sequences and Relationship Learning for Valence-Arousal Estimation

### 19 Mar 2024 v1

@@ Abstract
This paper presents our approach for the VA (Valence-Arousal) estimation task in the ABAW6 competition. We devised a comprehensive model by preprocessing video frames and audio segments to extract visual and audio features. Through the utilization of Temporal Convolutional Network (TCN) modules, we effectively captured the temporal and spatial correlations between these features. Subsequently, we employed a Transformer encoder structure to learn long-range dependencies, thereby enhancing the model’s performance and generalization ability. Our method leverages a multimodal data fusion approach, integrating pre-trained audio and video backbones for feature extraction, followed by TCN-based spatiotemporal encoding and Transformer-based temporal information capture. Experimental results demonstrate the effectiveness of our approach, achieving competitive performance in VA estimation on the AffWild2 dataset.
@@ 1 Introduction
In recent years, human emotional behavior analysis has received increasing attention. Commonly used human expression modalities in this field include Action Units (AU), basic expression categories (EXPR), and Valence-Arousal (VA). Action Units describe the movement of specific facial regions and are the smallest units for describing expressions. Basic expression categories classify expressions into several emotional categories such as happiness, sadness, etc. VA is a model that includes two continuous values representing the valence and arousal of emotions, which can better describe human emotional states.
The ABAW workshop aims to address the challenges of in-the-wild emotional behavior analysis, which reflects a key characteristic of real-world Human-Computer Interaction (HCI) systems. The objective is to develop intelligent machines and robots capable of understanding human emotions, moods, and behaviors. By interacting with humans in a ”human-centered” manner, these systems provide engaging experiences and effectively serve as digital assistants. Such interactions should not be constrained by individual backgrounds, age, gender, race, education level, occupation, or social status. Therefore, the development of intelligent systems capable of accurately analyzing human behavior in the wild is essential for fostering trust, understanding, and intimacy between humans and machines in real-life environments. The advancement of this technology holds promise for the future of HCI, enabling more sophisticated and human-centric experiences in everyday life. To promote the development of human emotional behavior analysis, the Affective Behavior Analysis in-the-wild (ABAW) competition has proposed a series of challenges aimed at overcoming obstacles in this field. The ABAW competition has constructed large-scale multimodal video datasets, such as Affwild [29, 14]and Affwild2[12, 13], which provide valuable resources for researchers, greatly advancing the progress of in-the-wild facial expression analysis and accelerating the application in related industries. The Affwild2 dataset contains a large number of videos, most of which are annotated frame by frame with AU, basic expression categories, VA, and other labels, providing researchers with rich annotated data.
@@
To address the challenges above, we employ a multimodal data fusion approach. Firstly, we utilize a pre-trained audio VGGnet as the backbone for extracting dynamic audio features from VGGish[9], and an IResnet[6] pre-trained with the arcface[5] method as the backbone for extracting video frames. Consequently, the expressiveness of deep features in both visual and audio modalities can be further enhanced through fine-tuning. Subsequently, for each branch, we utilize CNN backbones to extract dynamic spatial depth features from video frames and log-spectrograms, followed by the use of Temporal Convolutional Networks[1] to further learn spatiotemporal encoding. Lastly, these features are concatenated and fed into a Transformer structure to capture temporal information for downstream tasks. During training, we employ a strategy of large-window resampling, where the window size determines the amount of contextual information the model considers for predicting each time step.
@@
4. During training, we employ a large-window resampling strategy to determine the amount of contextual information considered for predicting each time step, enhancing the model’s ability to understand long-term dependencies in the data.
@@ 2 Related Work
Figure 1: Our proposed framework for VA estima-tion.
@@ 3.1 Feature Extraction
IResNet-50:IResNet-50[2] is a variant of the ResNet[8]architecture, which was introduced by Microsoft Research in their paper ”Deep Residual Learning for Image Recognition” by Kaiming He et al. IResNet-50 incorporates the idea of residual connections to enable the training of very deep neural networks more effectively. In this study, we trained IResNet-50 using the AffectNet[18] dataset and obtained a 512-dimensional visual feature vector.

### 20 Mar 2024 v2

@@ Abstract
This paper presents our approach for the VA (Valence-Arousal) estimation task in the ABAW6 competition. We devised a comprehensive model by preprocessing video frames and audio segments to extract visual and audio features. Through the utilization of Temporal Convolutional Network (TCN) module, we effectively captured the temporal and spatial correlations between these features. Subsequently, we employed a Transformer encoder structure to learn long-range dependencies, thereby enhancing the model’s performance and generalization ability. Additionally, we proposed the LA-SE module to better capture local image information and enhance channel selection and suppression. Our method leverages a multimodal data fusion approach, integrating pre-trained audio and video backbones for feature extraction, followed by TCN-based spatiotemporal encoding and Transformer-based temporal information capture. Experimental results demonstrate the effectiveness of our approach, achieving competitive performance in VA estimation on the AffWild2 dataset.
@@ 1 Introduction
In recent years, human emotional behavior analysis has received increasing attention. Commonly used human expression modalities[22, 20, 14] in this field include Action Units (AU)[23] basic expression categories (EXPR)[17], and Valence-Arousal (VA)[16]. Action Units describe the movement of specific facial regions and are the smallest units for describing expressions. Basic expression categories classify expressions into several emotional categories such as happiness, sadness, etc. VA is a model that includes two continuous values representing the valence and arousal of emotions, which can better describe human emotional states. The ABAW workshop[25, 24, 13] aims to address the challenges of in-the-wild emotional behavior analysis, which reflects a key characteristic of real-world Human-Computer Interaction (HCI) systems. The objective is to develop intelligent machines and robots capable of understanding human emotions, moods, and behaviors. By interacting with humans in a ”human-centered” manner, these systems provide engaging experiences and effectively serve as digital assistants. Such interactions should not be constrained by individual backgrounds, age, gender, race, education level, occupation, or social status. Therefore, the development of intelligent systems capable of accurately analyzing human behavior in the wild is essential for fostering trust, understanding, and intimacy between humans and machines in real-life environments. The advancement of this technology holds promise for the future of HCI, enabling more sophisticated and human-centric experiences in everyday life. To promote the development of human emotional behavior analysis, the Affective Behavior Analysis in-the-wild (ABAW) competition has proposed a series of challenges aimed at overcoming obstacles in this field. The ABAW competition has constructed large-scale multimodal video datasets, such as Affwild [41, 20, 21, 19]and Affwild2[15, 16, 13, 18], which provide valuable resources for researchers, greatly advancing the progress of in-the-wild facial expression analysis and accelerating the application in related industries. The Affwild2 dataset contains a large number of videos, most of which are annotated frame by frame with AU, basic expression categories, VA, and other labels, providing researchers with rich annotated data.
@@
To address the challenges above, we employ a multimodal data fusion approach. Firstly, we utilize a pre-trained audio VGGnet as the backbone for extracting dynamic audio features from VGGish[9], and an IResnet[6] pre-trained with the arcface[5] method as the backbone for extracting video frames. Consequently, the expressiveness of deep features in both visual and audio modalities can be further enhanced through fine-tuning. Subsequently, for each branch, we utilize CNN backbones to extract dynamic spatial depth features from video frames and log-spectrograms, followed by the use of Temporal Convolutional Networks[1] to further learn spatiotemporal encoding. Additionally, we integrate the LA-SE module before the TCN module to better capture local image information and enhance channel selection and suppression. The LA-SE module combines LANet for spatial aggregation and SENet for channel-wise selection, improving the effectiveness of feature extraction. Lastly, these features are concatenated and fed into a Transformer structure to capture temporal information for downstream tasks. During training, we employ a strategy of large-window resampling, where the window size determines the amount of contextual information the model considers for predicting each time step.
@@
2. To enhance feature extraction and selection, we propose the LA-SE network, combining LANet for spatial aggregation and SENet for channel-wise selection. This module captures local image information effectively, improving model performance.
@@ 2 Related Work
Figure 1: Our proposed framework for VA estimation. The model begins with preprocessing of video frames and audio segments, followed by feature extraction using pre-trained audio and video backbones. The LA-SE module enhances local image information capture and channel selection. Temporal Convolutional Network (TCN) modules capture temporal and spatial correlations in the features, while a Transformer encoder structure learns long-range dependencies.
@@ 3.1 Feature Extraction
IResNet-50:IResNet-50[2] is a variant of the ResNet[8] architecture. IResNet-50 incorporates the idea of residual connections to enable the training of very deep neural networks more effectively. In this study, we trained IResNet-50 using the AffectNet[29] dataset and obtained a 512-dimensional visual feature vector.
@@ 3.3 Training Details
3.4 LA-SE Module
To better capture local image information and enhance channel selection and suppression, we propose the LA-SE network, which combines the LANet[40] and SENet[10]. In Figure 2, LANet is employed to aggregate local information, accomplished through a sequence of two consecutive $1\times 1$ convolutional layers. These layers consolidate spatial information into a unified channel, which is then subjected to scaling using a sigmoid function to produce spatial attention.In contrast, Figure 3 illustrates the utilization of SENet for channel-wise selection and suppression. SENet comprises two primary operations: squeeze and excitation.
The squeeze operation compresses global channel information into a single-channel descriptor using global average pooling:
$z_{t}=\frac{1}{w\times h}\sum_{i=1}^{w}\sum_{j=1}^{h}u_{t}(i,j)$ (3)
where $z_{t}$ represents the signal of channel $t$ and $u_{t}(i,j)$ denotes the element of channel $t$ at position $(i,j)$. The excitation operation models channel-wise dependencies through two fully connected (FC) layers:
$s=\sigma\left(w_{2}\cdot g(w_{1}\cdot z)\right)$ (4)
where $\sigma$ refers to the sigmoid function, $g$ is the ReLU function, and $w_{1}$ and $w_{2}$ are the weights of the FC layers.
Finally, the scaling operation adjusts each channel dynamically based on learned activations:
$x_{i}=s_{i}\times u_{i}$ (5)
where $s_{i}$ represents a scalar related to channel $i$, and $u_{i}$ is the input for channel $i$.
The LA-SE network integrates the advantages of LANet and SENet to extract more representative and robust features for image understanding tasks, thereby improving model performance and generalization capability.
3.5 TCN Module
The Temporal Convolutional Network (TCN) module is utilized to capture temporal dependencies between frames and features. It consists of a stack of dilated convolutional layers followed by activation functions and pooling operations. The dilated convolutional layers have exponentially increasing dilation rates, enabling them to capture information from a wider range of temporal contexts.
Formally, given an input sequence $X=\{x_{1},x_{2},...,x_{T}\}$, where $x_{t}$ represents the feature representation at time step $t$, the output of the TCN module can be calculated as follows:
$y_{t}=\text{ReLU}\left(\sum_{i=1}^{N}W_{i}*x_{t+(i-1)d}\right)$ (6)
where *** denotes the convolution operation, $W_{i}$ represents the learnable weights of the $i$-th convolutional layer, $N$ is the number of layers, and $d$ is the dilation rate. The ReLU function introduces non-linearity to the model, allowing it to learn complex temporal patterns.
The output of the TCN module, denoted as $Y=\{y_{1},y_{2},...,y_{T}\}$, contains features that capture temporal dependencies and can be further processed by subsequent layers for downstream tasks such as classification or regression.
3.6 Transformer Encode
The Transformer Encoder is a crucial component in many sequence modeling tasks, including natural language processing and video analysis. It utilizes self-attention mechanisms to capture long-range dependencies within sequences effectively.
Formally, given an input sequence $\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{T}\}$, where $\mathbf{x}_{t}$ represents the feature representation at time step $t$, the Transformer Encoder processes the sequence through multiple layers of self-attention and feed-forward neural networks. Each layer consists of two sub-layers: a multi-head self-attention mechanism and a position-wise fully connected feed-forward network.
The multi-head self-attention mechanism computes attention scores between all pairs of positions in the input sequence, allowing the model to focus on relevant information from distant positions. This is achieved by linearly projecting the input into multiple lower-dimensional representations, computing attention scores, and then aggregating them across different heads.
The position-wise feed-forward network applies a two-layer fully connected neural network independently to each position in the sequence. This operation introduces non-linearity and enables the model to capture complex interactions between features.
After processing through multiple layers of self-attention and feed-forward networks, the Transformer Encoder produces a sequence of output representations $\mathbf{Y}=\{\mathbf{y}_{1},\mathbf{y}_{2},...,\mathbf{y}_{T}\}$, which captures rich contextual information and dependencies within the input sequence. These representations can be further used for downstream tasks such as classification, regression, or sequence generation.
Figure 2: The LANet module, where H, W and C refer to height, width and number of channels, respectively.
Figure 2: The LANet module, where H, W and C refer to height, width and number of channels, respectively.
Figure 3: The SENet module, where H, W and C refer to height, width and number of channels, respectively.
Table 2: Module Ablation Experiment Results for LA-SE, TCN, and Transformer Encode
Experimental Combination LA-SE TCN Transformer Encode CCC Score (%)
Baseline 31.08
+LA-SE 33.96
+TCN 51.57
+Transformer Encode 38.77
+LA-SE + TCN 52.71
+LA-SE + Transformer Encode 40.12
+TCN + Transformer Encode 60.21
+LA-SE + TCN + Transformer Encode 63.56
@@ 4.1 Dataset and Evaluation
4.2 Module Ablation Experiment Results
In this section, we present the results of the module ablation experiments for LA-SE, TCN, and Transformer Encode, along with their impact on the CCC (Concordance Correlation Coefficient) scores.
The table 2 presents the CCC scores (%) obtained from different experimental combinations. Each row represents a specific experimental configuration, including individual modules and their combinations. Starting with the Baseline model, which directly concatenates features extracted from video and audio, yielding a CCC score of 31.08. Subsequently, we gradually added different modules to investigate their effects on model performance. By adding the LA-SE module, the CCC score increased to 33.96, indicating an improvement in model performance. When the TCN module was added alone, the CCC score significantly increased to 51.57, demonstrating its effectiveness in capturing temporal dependencies between frames. The introduction of the Transformer Encode module further improved performance, increasing the CCC score to 38.77. This suggests that Transformer Encode helps to learn long-range dependencies between features, enhancing the model’s generalization capability. When combining the LA-SE and TCN modules, the CCC score further improved to 52.71, demonstrating the synergistic effect of these two modules. Similarly, the combination of LA-SE and Transformer Encode also led to a performance improvement, with a CCC score of 40.12. However, the most significant improvement occurred with the combination of TCN and Transformer Encode, resulting in a CCC score of 60.21. This indicates that TCN and Transformer Encode modules complement each other well in capturing temporal and feature relationships. Finally, when all three modules were combined, the model achieved the highest performance, with a CCC score of 63.71. This further validates the importance and effectiveness of the LA-SE, TCN, and Transformer Encode modules in improving model performance.